3. Find f“(3) if f(x) = V10x - 3 4. Find any if y = u² – 1 and u =
Answer #1

Please see the solution and If you have any doubt Please comment

Thanks

د) f (3) if f(x) = Niox-3 f(x) = (102-3) f(x) = 1 (10x - 3) - لری d (10x-3) da fice) = 1 th I (10x-3) 3 flick) = 10 3 a [10